Suppose that in a day, the probability of a car agent’'s not closing any deal is 0.35. on the other hand, the probability that he/she can close one deal is 0.3; two deals, 0.25; and three deals, 0.1. find the agent’s expected number of closed deals in a day and determine the variance and standard deviation.


1
Expert's answer
2022-04-12T17:14:27-0400

Let X - the random variable of the number of closed deals in a day.


The mean (the agent’s expected number of closed deals in a day):

"\mu=\sum x_i\cdot P(x_i)=\\
=0\cdot0.35+1\cdot0.3+2\cdot0.25+3\cdot0.1=1.1."


The variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\\
=\{ 0-1.1, 1-1.1, 2-1.1, 3-1.1\}=\\
=\{-1.1,-0.1,0.9,1.9\},"

"\sigma^2=(-1.1)^2\cdot0.35+(-0.1)^2\cdot0.3+0.9^2\cdot0.25+\\
+1.9^2\cdot0.1=0.99."


The standard deviation:

"\sigma=\sqrt{0.99}=0.995."